OpenStudy (dls):

find derivativeof e^xy+xy=0

OpenStudy (dls):

\[e^{xy} +xy=0\]

OpenStudy (anonymous):

what about it

OpenStudy (dls):

@zepdrix

OpenStudy (anonymous):

wrt to what?

OpenStudy (dls):

dy/dx

zepdrix (zepdrix):

:o

OpenStudy (dls):

O_o

zepdrix (zepdrix):

\[\huge e^{xy}(xy)'+(x)'y+x(y)'=0\]Which part you getting stuck on? :D Looks like we have the product rule a few times.

OpenStudy (dls):

I know e^x=e^x itself but e^xy too?

zepdrix (zepdrix):

\[\huge (e^{\text{potato}})'=e^{\text{potato}}(\text{potato})'\]Yup the thing in the exponent doesn't change :D You just need to apply the chain rule.

OpenStudy (dls):

okay! thanks :")